# Networked Communication for Decentralised Cooperative Agents in Mean-Field Control

### v1

@@ Abstract
We introduce networked communication to mean-field control (MFC) - the cooperative counterpart to mean-field games (MFGs) - and in particular to the setting where decentralised agents learn online from a single, non-episodic run of the empirical system. We adapt recent algorithms for MFGs to this new setting, as well as contributing a novel sub-routine allowing networked agents to estimate the global average reward from their local neighbourhood. We show that the networked communication scheme allows agents to increase social welfare faster than under both the centralised and independent architectures, by computing a population of potential updates in parallel and then propagating the highest-performing ones through the population, via a method that can also be seen as tackling the credit-assignment problem. We prove this new result theoretically and provide experiments that support it across numerous games, as well as exploring the empirical finding that smaller communication radii can benefit convergence in a specific class of game while still outperforming agents learning entirely independently. We provide numerous ablation studies and additional experiments on numbers of communication round and robustness to communication failures.
@@ 1 Introduction
Multi-agent reinforcement learning (MARL) can struggle to scale computationally as the number of agents N increases. The alternative mean-field game (MFG) framework (Lasry & Lions, 2007; Huang et al., 2006) has been used to address this scaling difficulty (Yardim & He, 2024; Zeng et al., 2024); it models a representative agent as interacting not with other individual agents in the population on a per-agent basis, but instead with a distribution over the other agents, known as the mean field. The framework analyses the limiting case when the population consists of an infinite number of symmetric and anonymous agents, that is, they have identical reward and transition functions which depend on the mean-field distribution rather than on the actions of specific other players. The MFG is a non-cooperative scenario where each agent seeks to maximise its individual return, and the solution to the game is a mean-field Nash equilibrium (MFNE), which can be used as an approximation for the Nash equilibrium (NE) in a finite-agent game, with the error in the solution reducing as N tends to infinity (Saldi et al., 2018; Anahtarci et al., 2023; Cui et al., 2023c; Yardim et al., 2024; Toumi et al., 2024; Hu & Zhang, 2024; Chen et al., 2024). Alternatively we can consider a cooperative scenario called a mean-field control (MFC) problem, where the population seeks to maximise a social welfare criterion such as the average return received by the agents. While MFGs have been well-studied and applied to a wide variety of real-world problems (Laurière et al., 2022a), MFC has received less attention, despite possibly being more useful for engineering collective behaviours to achieve global objectives (Cui et al., 2023c).
@@
Since MFC problems can be interpreted as optimisation problems from the perspective of a social planner, classical approaches to MFC involve centralised methods whereby a central learner updates a policy that is assumed to be passed automatically to the population (Fornasier & Solombrino, 2014; Carmona et al., 2019; Ruthotto et al., 2020; Laurière et al., 2022a; Angiuli et al., 2022; 2023; Cui et al., 2023a; Lee et al., 2024; Denkert et al., 2024). Often the empirical mean field of the actual population is not used, with the central learner updating an estimate of the mean field based on its own policy (Carmona et al., 2019; Angiuli et al., 2022; 2023). However, recent works on MFGs, as in other areas of multi-agent research, have recognised that the existence of a central coordinator is a strong assumption in complex, real-world settings, as well as representing a bottleneck for computation and communication, and a vulnerable single point of failure of the system (Wai et al., 2018; Zhang et al., 2018; 2021a; 2021b; Chen et al., 2021; Jiang et al., 2024; Xu et al., 2025; Agyeman et al., 2025; Cui et al., 2023c). These works also argue that, in order to be applicable to real-world problems such as swarm robotics, other desirable qualities for mean-field algorithms include: learning from the population’s empirical mean field (i.e. this distribution is generated only by the agents’ policies, rather than being updated by the algorithm itself or an external oracle/simulator); learning online from a single, non-episodic system run (i.e. similar to above, the population cannot be arbitrarily reset by an external controller); model-free learning; and function approximation to allow high-dimensional observations (Yardim et al., 2023; Benjamin & Abate, 2023; 2024).
@@
Some recent MFC works have considered decentralisation, but Bayraktar & Kara (2024) requires that decentralised agents optimise for learnt models of the system dynamics (and is only fully independent when the population is large but finite rather than infinite), while Cui et al. (2023c) presents a model-free deep learning algorithm that gives decentralised execution but requires centralised, episodic training. This latter work stipulates that decentralised training can be achieved if all agents can directly observe the mean-field distribution and use the same seed to correlate their actions (though they only provide empirical results for the centralised scenario, and Bayraktar & Kara (2024) provides no empirical results at all). However, assuming decentralised agents have access to this global information is unrealistic, and Benjamin & Abate (2024) in the non-cooperative MFG setting has shown that networked communication between decentralised agents allows agents to estimate the global mean field from a local neighbourhood. They also show that proliferating high-performing policies through the population via decentralised communication (in a manner reminiscent of distributed embodied evolutionary algorithms (Hart et al., 2015; Fernández Pérez et al., 2018; Fernández Pérez & Sanchez, 2019; Cazenille et al., 2025)) improves training time and avoidance of local optima, particularly over the case of agents learning entirely independently, but often also over populations with a single central learner.
@@
Inspired by this non-cooperative MFG work, we introduce networked communication to MFC for the first time, where populations arguably have even more incentive to communicate. This allows us to present a model-free deep learning algorithm that fulfils all of the proposed desiderata, including learning online from a single non-episodic run of the empirical system, and decentralised training without needing to observe global information: we contribute a novel sub-routine for estimating the global average reward from local communication, in addition to the existing sub-routine for estimating the global mean field from Benjamin & Abate (2024). We contribute theoretical proofs that decentralised policy communication allows networked populations to learn faster than both the independent and the centralised alternatives in the MFC setting in multiple classes of cooperative game. We also demonstrate this finding empirically in numerous games, as well as contributing an empirical study of the algorithms’ robustness to communication failures, along with several ablation studies. Please see the cover page for a summary of our contributions.
@@
The rest of this work is organised as follows. We give preliminaries in Sec. 2, and our learning and estimation algorithms in Sec. 3. We present theoretical results in Sec. 4, and experiments in Sec. 5. We give a more detailed comparison to related work in Appx. D.
@@ 2 Preliminaries
Solving the MFC problem involves finding the single policy that, when given to all agents in the infinite population, maximises the population’s expected return. We give two ways to conceive of our work, making more explicit the motivations underpinning other MFC works, illustrated in Fig. 2 (Appx. A) (Cui et al., 2023c; Dayanikli et al., 2024; Zaman et al., 2024; Bayraktar & Kara, 2024). Firstly, we contribute algorithms that allow the solution to a MFC problem to be learnt using the empirical distribution of a decentralised finite population, without needing to make unrealistic assumptions about access to an oracle for the infinite population. Note that empirically it may be impractical to assume that the decentralised agents always follow a single identical policy.
Alternatively, we may have originally been interested in solving a cooperative problem for a large, finite population, but, due to the scalability issues of learning approaches like MARL, forced to turn to the MFC framework to find a policy that gives an approximate solution to the finite-population problem. We contribute algorithms that allow the deployed finite population to find the MFC solution that in turn approximately solves the original problem, without unrealistic assumptions about centralised training. Under this framing, it may matter less whether all agents follow a single policy in practice (Yardim et al. (2023); Benjamin & Abate (2023; 2024) follow a similar logic in MFGs).
@@ 2.1 Mean-field control
We use the following notation. $N$ is the number of agents in a population, with $\mathcal{S}$ and $\mathcal{A}$ representing the finite state and common action spaces. The set of probability measures on a finite set $\mathcal{X}$ is denoted $\Delta_{\mathcal{X}}$, and $\mathbf{e}_{x}\in\Delta_{\mathcal{X}}$ for $x\in\mathcal{X}$ is a one-hot vector with only the entry corresponding to $x$ set to 1, and all others set to 0. For time $t\geq 0$, $\hat{\mu}_{t}$ = $\frac{1}{N}\sum^{N}_{i=1}\sum_{s\in\mathcal{S}}$ $\mathds{1}_{s^{i}_{t}=s}\mathbf{e}_{s}$in $\Delta_{\mathcal{S}}$ is a vector of length $|\mathcal{S}|$ denoting the empirical categorical state distribution of the $N$ agents at time $t$. For agent $i\in\{1\dots N\}$, $i$’s policy $\pi^{i}\in\Pi$ depends on its observation $o^{i}_{t}$. We give different forms that this observation can take, and relatedly a more formal definition of the policy, after the following.
Definition 1 (N-player stochastic cooperative control problem with symmetric, anonymous agents).
@@
$s^{i}_{t+1}\sim P(\cdot|s^{i}_{t},a^{i}_{t},\hat{\mu}_{t}),\;\;\;r^{i}_{t}=R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t}).$
For the joint policy $\boldsymbol{\pi}$ := ($\pi^{1},\dots,\pi^{N}$) $\in\Pi^{N}$, an individual agent’s discounted return is given by:
Definition 2 (Individual expected discounted return).
For all $i,j\in\{1,\dots,N\}$, $i$’s return is
$V^{i}(\boldsymbol{\pi},\mu_{0})=\mathbb{E}\left[\sum^{\infty}_{t=0}\gamma^{t}R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})\bigg{|}\begin{subarray}{c}s^{j}_{0}\sim\mu_{0}\\ a^{j}_{t}\sim\pi^{j}(o^{j}_{t})\\ s^{j}_{t+1}\sim P(\cdot|s^{j}_{t},a^{j}_{t},\hat{\mu}_{t})\end{subarray}\right].$
@@
Definition 3 (Population-average expected discounted return).
@@
$V^{pop}(\boldsymbol{\pi},\mu_{0})=\frac{1}{N}\sum_{i}^{N}V^{i}(\boldsymbol{\pi},\mu_{0})=\mathbb{E}\left[\frac{1}{N}\sum^{\infty}_{t=0}\sum_{i}^{N}\gamma^{t}R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})\bigg{|}\begin{subarray}{c}s^{j}_{0}\sim\mu_{0}\\ a^{j}_{t}\sim\pi^{j}(o^{j}_{t})\\ s^{j}_{t+1}\sim P(\cdot|s^{j}_{t},a^{j}_{t},\hat{\mu}_{t})\end{subarray}\right].$
That is, the solution to the control problem is $\boldsymbol{\pi}^{*}=\arg\max_{\boldsymbol{\pi}\in\Pi^{N}}V^{pop}(\boldsymbol{\pi},\mu_{0})$.
@@
At the limit as $N\rightarrow\infty$, the infinite population of agents can be characterised as a limit distribution $\mu\in\Delta_{\mathcal{S}}$; the infinite-agent setting is termed a MFC problem. The so-called ‘mean-field flow’ $\boldsymbol{\mu}$ is given by the infinite sequence of mean-field distributions s.t. $\boldsymbol{\mu}=(\mu_{t})_{t\geq 0}$.
Definition 4 (Induced mean-field flow).
@@
We denote by $I(\pi)$ the mean-field flow $\boldsymbol{\mu}$ induced when all the agents follow $\pi$, where this is generated from $\pi$ by $\mu_{t+1}(s^{\prime})=\sum_{s,a}\mu_{t}(s)\pi(a|o_{t})P(s^{\prime}|s,a,\mu_{t})$. The snapshot of this induced flow at $t$ is given by $I(\pi)_{t}$.
Definition 5 (Social welfare).
@@
When all agents follow policy $\pi$ giving mean-field flow $\boldsymbol{\mu}=I(\pi)$, the social welfare of the policy is given by
@@
$W(\pi;I(\pi))=\mathbb{E}\left[\sum^{\infty}_{t=0}\gamma^{t}(R(s_{t},a_{t},I(\pi)_{t}))\bigg{|}\begin{subarray}{c}s_{0}\sim\mu_{0}\\ a_{t}\sim\pi(\cdot|o_{t})\\ s_{t+1}\sim P(\cdot|s_{t},a_{t},{I(\pi)_{t}})\end{subarray}\right].$
Definition 6 (Social optimum).
@@
The solution to the MFC problem is a social optimum policy $\pi^{*}\in\Pi$ that maximises the social welfare function in Def. 5, i.e. $\pi^{*}=\arg\max_{\pi\in\Pi}W(\pi;I(\pi))$.
Remark 1.
[…]
@@
When the distribution is the same for all $t$, i.e. $\mu_{t}=\mu_{t+1}$ $\forall t\geq 0$, we say the mean-field flow is stationary, giving a stationary MFC problem. Non-stationary problems require the policy to depend on the mean field such that $o^{i}_{t}=(s^{i}_{t},\hat{\mu}_{t})$, whereas the observation in the stationary case can be simplified to $o^{i}_{t}=s^{i}_{t}$. However, classical approaches to the MFC problem that conceive of a central planner trying to guide the population to a distribution that maximises the expected return might have policies that depend on the mean field even in the stationary case (Laurière et al., 2022a; Carmona et al., 2023; Cui et al., 2023c). Therefore we use mean field-dependent policies for the sake of generality, but show through our ablation studies that in practice our algorithms require only $\pi^{i}(a|o^{i}_{t})=\pi^{i}(a|s^{i}_{t})$ in our experimental tasks, which have stationary solutions.
@@
Furthermore, it is unrealistic to assume that decentralised agents with a possibly limited communication radius would have perfect observability of the global mean field $\hat{\mu}_{t}$. Therefore we allow agents to form a local estimate $\tilde{\hat{\mu}}^{i}_{t}$ which can be improved by communication with neighbours, using Alg. 3 (from Alg. 3 in Benjamin & Abate (2024) for the MFG setting). We thus have $o^{i}_{t}=(s^{i}_{t},\tilde{\hat{\mu}}^{i}_{t})$. Formally we can now say that when $o^{i}_{t}=(s^{i}_{t},\hat{\mu}_{t})$ or $(s^{i}_{t},\tilde{\hat{\mu}}^{i}_{t})$, we have the set of policies defined as $\Pi$ = $\pi$ : $\mathcal{S}\times\Delta_{\mathcal{S}}\rightarrow\Delta_{\mathcal{A}}$}, and the set of Q-functions denoted $\mathcal{Q}=\{q:\mathcal{S}\times\Delta_{\mathcal{S}}\times\mathcal{A}\rightarrow\mathbb{R}\}$. [^1] When $o^{i}_{t}=s^{i}_{t}$, we instead have $\Pi$ = $\pi$ : $\mathcal{S}\rightarrow\Delta_{\mathcal{A}}$} and $\mathcal{Q}=\{q:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}\}$.

### v2

@@ Abstract
The mean-field framework has been used to find approximate solutions to problems involving very large populations of symmetric, anonymous agents, which may be intractable by other methods. The cooperative mean-field control (MFC) problem has received less attention than the non-cooperative mean-field game (MFG), despite the former potentially being more useful as a tool for engineering large-scale collective behaviours. Decentralised communication algorithms have recently been introduced to MFGs, giving benefits to learning speed and robustness. Inspired by this, we introduce networked communication to MFC - where populations arguably have broader incentive to communicate - and in particular to the setting where decentralised agents learn online from a single, non-episodic run of the empirical system. We adapt recent MFG algorithms to this new setting, as well as contributing a novel sub-routine allowing networked agents to estimate the global average reward from their local neighbourhood. Previous theoretical analysis of decentralised communication in MFGs does not extend trivially to MFC. We therefore contribute new theory proving that in MFC the networked communication scheme allows agents to increase social welfare faster than under both of the two typical alternative architectures, namely independent and centralised learning. We provide experiments that support this new result across different classes of cooperative game, and also give numerous ablation studies and additional experiments concerning numbers of communication round and robustness to communication failures.
@@ 1 Introduction
The mean-field framework (lasry2007mean; huangMFG) models a representative agent as interacting not with the rest of the population on a per-agent basis, but instead with a distribution over the other agents, known as the mean field. The framework analyses the limiting case when the population consists of an infinite number of symmetric and anonymous agents, that is, they have identical reward and transition functions which depend on the mean-field distribution rather than on the actions of specific other players. The mean-field control (MFC) problem is a cooperative scenario where the population seeks to maximise a social welfare criterion such as the average return received by the agents. Alternatively we can consider a non-cooperative scenario called a mean-field game (MFG), where each agent seeks to maximise its individual return, to which the solution is a MFG-Nash equilibrium (MFG-NE).
The MFC social optimum and the MFG-NE can respectively be used as approximate solutions to the associated finite-agent problem/game, with the error in the solution reducing as the number of agents N tends to infinity (finite_proofs; doi:10.1137/20M1360700; 10.5555/3586589.3586718; major_minor_control; cui2023learningdecentralisedcontrol; Anahtarci2020QLearningIR; yardim2024mean; TOUMI2024111420; hu2024mfoml; chen2024periodic; bayraktar2024learning). MFC and MFGs have therefore been used to address the difficulty faced by multi-agent reinforcement learning (MARL), which can struggle to scale computationally as N increases (yardim2024exploiting; zeng2024single). While MFGs have been well-studied and applied to a wide variety of real-world problems (survey_learningMFGs), MFC has received less attention, despite possibly being more useful for engineering collective behaviours to achieve global objectives, such as in consensus, synchronisation, rendezvous, exploration, coverage or task allocation problems (cui2023learningdecentralisedcontrol). This paper seek to redress some of this imbalance by adapting recent developments in MFGs for MFC.
Since MFC problems can be interpreted as optimisation problems from the perspective of a social planner, classical approaches involve centralised methods (they also do so for reasons of simplicity, as in MFGs). In this context ‘centralised’ does not necessarily imply global observability of the whole population’s actions - which could make computation infeasible given the complexity of the problem - but rather that learning is conducted from the samples of a single representative agent, whose policy updates are assumed to be automatically pushed to the rest of the population by the central node (Fornasier_Solombrino_2014; carmona2019linear; doi:10.1073/pnas.1922204117; survey_learningMFGs; angiuli2021unified; angiuli2023convergence; cui2023multiagent; lee2024mean; denkert2024randomisation). For this reason, whilst ‘centralised learning’ is the term used in prior works, we generally refer to ‘central-agent learning’ to reduce confusion. Often the empirical mean field of the actual population is not even used to compute rewards or transitions, with the central learner instead updating an estimate of the mean field based only on its own policy, which is in turn used as input to its reward and transition functions (carmona2019linear; angiuli2021unified; angiuli2023convergence).
@@
Recent works on MFGs, as in other areas of multi-agent research, have recognised that the existence of a central learner is a strong assumption in complex, real-world settings, as well as representing a bottleneck for computation and communication, and a vulnerable single point of failure of the system (mainone; decentralised_review; marlreview; distributed_review; policy_mirror_independent; benjamin2024networked; benjamin2024networkedapproximation; jiang2024fully; xu2025targets; AGYEMAN2025106183; 10971233). They advocate instead for the individual agents in the empirical population to learn policies for themselves without relying on a central node. Such works also argue that other strong classical assumptions should similarly be loosened in order to make MFGs applicable to real-world, embodied problems such as swarm robotics. They therefore contend that, aside from decentralised learning, desirable qualities for mean-field algorithms include: learning from the population’s empirical mean field (i.e. the mean-field distribution is generated only by the agents’ policies, rather than being manipulated by the algorithm itself or by an external oracle/simulator); learning online from a single, non-episodic system run (i.e. similar to above, the population is not arbitrarily reset by an external controller); learning without reliance on a model of the system; and using function approximation to allow scalability to high-dimensional observations (including the option to include the mean field in the input to policies).
@@
Until now, no work on MFC has met all these criteria. Some recent works have considered decentralisation in MFC, but bayraktar2024learning requires that decentralised agents optimise for learnt models of the system dynamics (and learning is only fully independent when the population is large but finite rather than infinite), while cui2023learningdecentralisedcontrol presents a model-free deep learning algorithm that gives decentralised execution but requires centralised, episodic training. This latter work stipulates that decentralised training can be achieved if all agents can directly observe the mean-field distribution and use the same seed to correlate their actions, though they only provide empirical results for the centralised scenario, while bayraktar2024learning provides no empirical results at all. However, assuming decentralised agents have access to this global information is unrealistic, and in the non-cooperative MFG setting benjamin2024networkedapproximation have shown that networked communication between decentralised agents allows agents to estimate the global mean field from a local neighbourhood. They also show that proliferating high-performing policies through the population via decentralised communication (in a manner reminiscent of distributed embodied evolutionary algorithms (survivability; Maintaining_Diversity; 10.1007/978-3-030-16692-2_38; cazenille2025signalling; sissodia2025evolutionary)) improves training time and avoidance of local optima, particularly over the case of agents learning entirely independently, but often also over populations with a single central learner.
@@
Inspired by this non-cooperative MFG work, we introduce networked communication to MFC for the first time, where populations arguably have even more incentive to communicate. This allows us to present a model-free deep learning algorithm that fulfils all of the proposed desiderata, including learning online from a single non-episodic run of the empirical system, and decentralised training without needing to observe global information: we contribute a novel sub-routine for estimating the global average reward from local communication, in addition to the existing sub-routine for estimating the global mean field from benjamin2024networkedapproximation. Previous theoretical analysis of networked communication in the non-cooperative MFG setting does not extend trivially to MFC, so we contribute new theoretical proofs showing that decentralised policy exchange allows networked populations to learn faster than both the independent and the central-agent alternatives in the MFC setting, across different classes of cooperative game (coordination and anti-coordination). We also demonstrate this finding empirically in numerous games, as well as contributing an empirical study of the algorithms’ robustness to communication failures, along with several ablation studies. In summary, our contributions include:
  • We provide the first algorithms in MFC for model-free training without any central provision of information or coordination, as well as the first MFC algorithms for online learning from a single, non-episodic run of the empirical system.
    • We contribute a novel sub-routine allowing decentralised agents to estimate the global average reward via networked communication, and incorporate an existing sub-routine used in MFGs for estimating the global mean field via local communication.
  • We prove theoretically that in this context, decentralised networked communication can improve learning speed over the independent and central-agent architectures.
  • We provide extensive experiments supporting our theoretical results in numerous games, and give ablation studies of various parts of our algorithms, as well as a study of robustness to communication failures.
@@
We provide further comparison with related work in Sec. 2, give preliminaries in Sec. 3, and our algorithms in Sec. 4. We present theoretical results in Sec. 5 and experiments in Sec. 6, before suggesting future work in Sec. 7.
2 Related work
We discuss here the research most closely related to our present work, focusing on decentralisation and networked communication, and clarifying the differences with prior methods and settings. We refer the reader to survey_learningMFGs for a broader survey of MFC.
Numerous works claiming to study decentralisation in MFC take this to mean only that agents do not have access to the specific states of all other agents, and have policies depending on their local state and possibly the mean field, all of which we take as a given in our work. They nevertheless rely on a central learner or coordinator that provides global information to all agents, a dependence that we remove in our work. This applies, for example, to 7368131, where a ‘central population coordinator’ broadcasts a common signal to all agents, and to tajeddini2017robust, which presents a leader-follower setting where a ‘central population coordinator’ estimates the mean-field trajectory. 8879544 similarly requires a central coordinator, and also presents a non-cooperative scenario so does not actually fall under MFC despite being referred to as such.
In cui2023learningdecentralisedcontrol, decentralisation applies only during execution, and they offer a centralised-training decentralised-execution method (as also in cui2023multiagent). They say that decentralised training could be achieved if the global mean field is observable and all agents use the same seed to correlate their actions, whilst we do not require either assumption for our decentralised training algorithm. They also train episodically whereas we learn online from a single run of the system. Finally, their experiments focus only on coordination games, whereas we additionally explore empirical effects resulting from decentralised training in anti-coordination games, where agents can gain higher rewards by diversifying their behaviour.
bayraktar2024learning considers independent, ‘online’ learning for MFC in a setting that is different from ours. Crucially, their method involves agents first estimating a model (reward and transition functions) of the system by conducting ‘online’ updates using samples collected while following exploration policies. Only once having done so do they compute execution policies that are optimal with respect to the estimated model. We argue that having a dedicated exploration phase is infeasible for many real-world applications, and instead present a fully model-free online learning algorithm. Moreover, their setting only permits independent learning if $N$ is large but finite. For infinite populations, a central coordinator is required to supply common noise to aid exploration during the initial phase, and if the optimal policy for the estimated model is not unique, centralised coordination is required to allow the agents to agree on which policy to execute. Our algorithm requires no such special considerations. Finally, their work is purely theoretical, whereas we provide extensive empirical results.
angiuli2021unified and angiuli2023convergence provide algorithms for MFC learning from a single run, but there it is a single run only of a ‘representative’ player that is used to simulate the mean field, rather than a single run of the empirical population as in our work. Their algorithms are thus inherently centralised, as well as involving two timescales for updating the mean-field approximation, which we argue is unlikely to be a practical paradigm for training in complex real-world systems such as robotic swarms.
Our work is also closely related to benjamin2024networked and benjamin2024networkedapproximation, which introduce networked communication to the non-cooperative MFG setting. By adapting their communication scheme and learning algorithm, we introduce networked communication to the cooperative MFC setting, where it is arguably more applicable due to broader incentives for agents to communicate policies. Their works focus on coordination games to justify the sharing of policies (though benjamin2024networkedapproximation does demonstrate empirically that networked agents outperform independent agents in a non-cooperative anti-coordination game, indicating that self-interested agents do nevertheless have incentive to communicate), whilst we provide extensive theoretical and empirical results on the benefits of policy sharing in MFC for both coordination and anti-coordination games. We leverage Alg. 4 from benjamin2024networkedapproximation for estimating the global mean field from a local neighbourhood, but additionally contribute the novel Alg. 1 for estimating the global average reward from a local neighbourhood for the MFC setting.
@@ 3.1 Mean-field control
We use the following notation. $N$ is the number of agents in a population, with $\mathcal{S}$ and $\mathcal{A}$ representing the finite state and common action spaces. The set of probability measures on a finite set $\mathcal{X}$ is denoted $\Delta_{\mathcal{X}}$, and $\mathbf{e}_{x}\in\Delta_{\mathcal{X}}$ for $x\in\mathcal{X}$ is a one-hot vector with only the entry corresponding to $x$ set to 1, and all others set to 0. For time $t\geq 0$, $\hat{\mu}_{t}$ = $\frac{1}{N}\sum^{N}_{i=1}\sum_{s\in\mathcal{S}}$ $\mathds{1}_{s^{i}_{t}=s}\mathbf{e}_{s}$ $\in$ $\Delta_{\mathcal{S}}$ is a vector of length $|\mathcal{S}|$ denoting the empirical categorical state distribution of the $N$ agents at time $t$. For agent $\in\{1\dots N\}, ii$’s policy $\pi^{i}\in\Pi$ depends on its observation $o^{i}_{t}$. We give different forms that this observation can take, and relatedly a more formal definition of the policy, after the following.
Definition 3.1 (N-player stochastic cooperative control problem with symmetric, anonymous agents).
@@
For the joint policy $\bm{\pi}$ := ($\pi^{1},\dots,\pi^{N}$) $\in\Pi^{N}$, an individual agent’s discounted return is given by:
Definition 3.2 (Individual expected discounted return).
For all $j\in\{1,\dots,N\}, ii$’s return is
$V^{i}(\bm{\pi},\mu_{\bar{t}})=\mathbb{E}\left[\sum^{\infty}_{t={\bar{t}}}\gamma^{t}R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})\bigg|\begin{subarray}{c}s^{j}_{\bar{t}}\sim\mu_{\bar{t}}\\ a^{j}_{t}\sim\pi^{j}(o^{j}_{t})\\ s^{j}_{t+1}\sim P(\cdot|s^{j}_{t},a^{j}_{t},\hat{\mu}_{t})\end{subarray}\right].$
@@
Definition 3.3 (Population-average expected discounted return).
@@
$V^{pop}(\bm{\pi},\mu_{\bar{t}})=\frac{1}{N}\sum_{i}^{N}V^{i}(\bm{\pi},\mu_{\bar{t}})=\mathbb{E}\left[\frac{1}{N}\sum^{\infty}_{t={\bar{t}}}\sum_{i}^{N}\gamma^{t}R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})\bigg|\begin{subarray}{c}s^{j}_{\bar{t}}\sim\mu_{\bar{t}}\\ a^{j}_{t}\sim\pi^{j}(o^{j}_{t})\\ s^{j}_{t+1}\sim P(\cdot|s^{j}_{t},a^{j}_{t},\hat{\mu}_{t})\end{subarray}\right].$
That is, the solution to the control problem is $\bm{\pi}^{*}=\arg\max_{\bm{\pi}\in\Pi^{N}}V^{pop}(\bm{\pi},\mu_{\bar{t}})$.
@@
At the limit as $N\rightarrow\infty$, the infinite population of agents can be characterised as a limit distribution $\mu\in\Delta_{\mathcal{S}}$; the infinite-agent setting is termed a MFC problem. The mean-field flow $\bm{\mu}$ is given by the infinite sequence of mean-field distributions s.t. $\bm{\mu}=(\mu_{t})_{t\geq 0}$.
Definition 3.4 (Induced mean-field flow).
@@
We denote by $I(\pi)$ the mean-field flow $\bm{\mu}$ induced when all the agents follow $\pi$, where this is generated from $\pi$ by
$\mu_{t+1}(s^{\prime})=\sum_{s,a}\mu_{t}(s)\pi(a|o_{t})P(s^{\prime}|s,a,\mu_{t}).$
The snapshot of this induced flow at $t$ is given by $I(\pi)_{t}$.
Definition 3.5 (Social welfare).
@@
When all agents follow policy $\pi$ giving mean-field flow $\bm{\mu}=I(\pi)$, $\pi$’s social welfare is
@@
$W(\pi;I(\pi))=\mathbb{E}\left[\sum^{\infty}_{t={\bar{t}}}\gamma^{t}(R(s_{t},a_{t},I(\pi)_{t}))\bigg|\begin{subarray}{c}s_{\bar{t}}\sim\mu_{\bar{t}}\\ a_{t}\sim\pi(\cdot|o_{t})\\ s_{t+1}\sim P(\cdot|s_{t},a_{t},{I(\pi)_{t}})\end{subarray}\right].$
Definition 3.6 (Social optimum).
@@
The solution to the MFC problem is a social optimum policy $\pi^{*}\in\Pi$ that maximises the social welfare function in Def. 3.5, i.e. $\pi^{*}=\arg\max_{\pi\in\Pi}W(\pi;I(\pi))$.
Remark 3.7.
Previous works showed that the MFC social optimum $\pi^{*}$ gives a good approximation for the harder-to-solve finite-agent problem (i.e. if $\bm{\pi}=(\pi^{*},\dots,\pi^{*})$), with the error characterised by $\mathcal{O}(\frac{1}{\sqrt{N}})$ (doi:10.1137/20M1360700; 10.5555/3586589.3586718; major_minor_control; cui2023learningdecentralisedcontrol; bayraktar2024learning).
@@
When the distribution is the same for all $t$, i.e. $\mu_{t}=\mu_{t+1}$ $\forall t\geq 0$, we say the mean-field flow is stationary, giving a stationary MFC problem. Non-stationary problems may require the policy to depend on the mean field such that $o^{i}_{t}=(s^{i}_{t},\hat{\mu}_{t})$, whereas the observation in the stationary case can be simplified to $o^{i}_{t}=s^{i}_{t}$. However, since classical approaches to the MFC problem often conceive of a central planner trying to guide the population to a distribution that maximises the expected return, they sometimes have policies that depend on the mean field even in the stationary case (survey_learningMFGs; carmona2021modelfree; cui2023learningdecentralisedcontrol). Therefore we permit mean field-dependent policies for the sake of generality, but show through our ablation studies that in practice our algorithms require only $\pi^{i}(a|o^{i}_{t})=\pi^{i}(a|s^{i}_{t})$ in our experimental tasks, which have stationary solutions.
@@
Furthermore, it is unrealistic to assume that decentralised agents with a possibly limited communication radius would have perfect observability of the global mean field $\hat{\mu}_{t}$. Therefore we allow agents to form a local estimate $\tilde{\hat{\mu}}^{i}_{t}$ which can be improved by communication with neighbours, using Alg. 4 (from Alg. 3 in benjamin2024networkedapproximation for the MFG setting). We thus have $o^{i}_{t}=(s^{i}_{t},\tilde{\hat{\mu}}^{i}_{t})$. Formally we can now say that when $o^{i}_{t}=(s^{i}_{t},\hat{\mu}_{t})$ or $(s^{i}_{t},\tilde{\hat{\mu}}^{i}_{t})$, we have the set of policies defined as $\Pi$ = $\pi$ : $\mathcal{S}\times\Delta_{\mathcal{S}}\rightarrow\Delta_{\mathcal{A}}$}, and the set of Q-functions denoted $\mathcal{Q}=\{q:\mathcal{S}\times\Delta_{\mathcal{S}}\times\mathcal{A}\rightarrow\mathbb{R}\}$. (N.b. when $o^{i}_{t}=s^{i}_{t}$, we instead have $\Pi$ = $\pi$ : $\mathcal{S}\rightarrow\Delta_{\mathcal{A}}$} and $\mathcal{Q}=\{q:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}\}$.)
Figure 1: Two possible ways to conceive of our work regarding the relationship between the infinite- and finite-population control problems, described in Rem. 3.8. Note that using the finite empirical population to learn the single-policy MFC social optimum $\bm{\pi}=(\pi^{*},\dots,\pi^{*})$ for the infinite population (Def. 3.6) is not the same as directly finding $\bm{\pi}^{*}=\arg\max_{\bm{\pi}\in\Pi^{N}}V^{pop}(\bm{\pi},\mu_{\bar{t}})=(\pi^{1},\dots,\pi^{N})$, i.e. the tuple of individual policies that maximises the expected finite population-average return in Def. 3.3, a problem known to be hard (cui2023learningdecentralisedcontrol; b57d2987-f484-384d-ab52-5bac787e11f2).
Remark 3.8.
We have just seen that an optimal solution to the theoretical MFC problem is a single policy that, when followed by all agents in the infinite population, maximises the population’s expected return. We give two ways to conceive of our work, illustrated in Fig. 1, which mirror and make more explicit the motivations underpinning other MFC works (cui2023learningdecentralisedcontrol; dayanikli2024deep; zaman2024robust; bayraktar2024learning; 10928821).
  1. 1.
    Firstly, while previous works might make unrealistic assumptions about access to an oracle for the infinite population, we contribute algorithms that allow the solution to a MFC problem to be learnt using the empirical distribution of a decentralised finite population. Note that it is unnecessary (and may be impractical) to assume the decentralised agents always follow a single identical policy throughout training.
  2. 2.
    Alternatively, we may have originally been interested in solving a cooperative problem for a large, finite population, but, due to the scalability issues of learning approaches like MARL, were forced to turn to the MFC framework to find a policy that gives an approximate solution to the finite-population problem. We contribute algorithms that allow the deployed finite population to find the MFC solution that in turn approximately solves the original problem, without unrealistic assumptions about centralised training. Under this framing, it may matter less whether all agents follow a single policy in practice (policy_mirror_independent and benjamin2024networked; benjamin2024networkedapproximation follow a similar logic in MFGs).